Coves Cakes is a local bakery. Price and cost information follows: $ 17.00 points Price per cake Variable cost per cake IngrCoves Cakes is a local bakery. Price and cost information follows: $ 17.00 points Price per cake Variable cost per cake IngrCoves Cakes is a local bakery. Price and cost information follows: $ 17.00 points Price per cake Variable cost per cake Ingr
Answer #1

1) Break even unit = Fixed cost/Contribution margin per unit = 3850/(17-4.10) = 298 Units

Break even sales = 298*17 = 5066

2) Margin of safety = (450-298)*17 = 2584

3) Target sales = (3850+2000)/12.90 = 453 Units